# Automatic Curriculum Design for Zero-Shot Human-AI Coordination

### v1

@@ 3.3 Learning Potential Score Based on Return
Our method follows the replay distribution (1) when selecting an environment for the ego-agent from the individual co-player buffer. $P_{S}$ is the first prioritization distribution in the replay distribution (1) and it assigns higher priority to environments with learning potential to train the ego-agent. Return is the appropriate scoring metric to construct a prioritization distribution in human-AI coordination. It represents the coordination performance during the episode and indicates which environment requires more training. If the replayed environment has a low return, it indicates that the ego-agent is hard to coordinate in that environment. The details of $S$ are as follows :
@@
Equation (3) is the expected return of the replayed environment with the sampled co-player during the episode. Here, $T$ represents the total number of timesteps in the episode, and $i$ denotes the index of the replayed environment from the co-player’s buffer. This environment score is normalized using Equation (4). We normalized the scores using a rank-based prioritization which sorts the scores in the co-player buffer in ascending order prioritizing the low-return environment. $j$ denotes the index of environments in the co-player’s buffer. $\beta$in [0, 1] is the range of values for the temperature parameter that adjusts the influence of a distribution.
@@ 4 Experiments Setup
Figure 3: (a) Overall results with a human proxy model. Our method achieves higher coordinate performance with the human proxy agent than other baselines on all five evaluation layouts (Mean and standard error were computed over 100 runs with three random seeds). (b) Utility function analysis ($Regret$ vs $Return$). Using a return-based utility function has an advantage to train ego-agent than the regret-base utility function in multi-agent coordination settings (Mean and standard error were computed over 100 runs with three random seeds).
@@
Figure 4: Learning curves for the training, training-evaluation and evaluation phases. The solid line shows the average episode reward for the different three random seeds, while the shaded areas show the maximum and minimum rewards. The results show that our model achieves higher performance in both the training and evaluation layouts.
Overcooked-AI Carroll et al. (2019) is a cooperative game played in a kitchen-inspired grid world environment. The main goal is to cook as much food as possible in a limited time. Agents can get rewards through completing the task sequentially such as ingredient preparation, cooking, and serving on plates. Cooperation is essential in this environment, where success or failure depends on the effectiveness of player coordination with their partner.
Baseline Model. We compare our method against three key baseline models related to UED: MAESTRO Samvelyan et al. (2023), Robust PLR Jiang et al. (2021b), and Domain Randomization Jakobi (1997). Robust PLR and domain randomization are methods originally used in a single-agent setting, we modified them for the multi-agent setting using the self-play method. The ego-agent was trained using the PPO method Schulman et al. (2017), which is well-known for training reinforcement learning agents. The model architecture and hyperparameter choices are in Appendix A.
Layout Generation. In our experiments, we automatically generated a total of 6,000 different layouts for Overcooked-AI, each layout size was 7 × 5. These layouts were generated using a heuristic-driven layout generator capable of adjusting three layout parameters: the number of interactive blocks (such as pots, plates, outlets, and onions), the number of spaces, and the size of layouts, which is fixed in this study. The generator was designed to ensure that each layout included at least 12 spaces and 5 interactive blocks. To prevent redundancy, the layout generator compared the layout array during generation. Additionally, check the layout solvability using the A* path-finding algorithm Hart et al. (1968), which can find the path for agent access to any interactive block. Further details on the algorithm are provided in Appendix B.
Human Proxy. To evaluate the coordination performance with the trained agent in our experiments, we utilized a human proxy agent from a previous study Yang et al. (2022) as a co-player agent. This human proxy agent performs as human-like as possible by adjusting its parameters based on human data using the cross-entropy methods.
Evaluation. We evaluated the human-AI zero-shot coordination performance on five unseen layouts (Figure 2), selected based on coordination difficulty. This helps to measure the adaptability of the ego-agent based on the difficulty of the layout. In addition, we conducted a utility function analysis.
@@ 5.1 Experiments With Human Proxy
Overall Result. Figure 3(a) shows the comparison result of our method and the baseline models (MAESTRO, Robust PLR, and Domain Randomization) regarding the coordination performance with a human proxy agent on unseen layouts. The bar graph is the evaluation result, which is the average episode reward (three random seeds) with the human proxy agent on the evaluation layouts. As a result, our method showed a better performance on all five evaluation layouts compared to other baselines. We observe that our method achieves a higher generalization coordination performance with unseen partners in unseen evaluation layouts. Furthermore, we also conducted cross-play experiments, where our method outperformed other baselines in unseen evaluation layouts. The results of these experiments are provided in Appendix C.
(a) Human Study Result.
(b) Human Survey Result.
Figure 5: This plot shows the results of an experiment with 20 real humans: (a) The average reward obtained from the five unseen evaluation layouts (b) The average evaluation scores for the two items "Collaborativeness" and "Human Preference". Each subject was scored on a scale of 1 to 4 based on their voting ranking. A higher score indicates a more positive evaluation.
@@ 5.2 Study of the Utility Function
Evaluation Result. We had an analysis to try to figure out which utility functions are appropriate for a Human-AI coordination setting. Figure 3(b) shows how the utility function affects the training of the ego-agent. Our method and Robust PLR improved coordination performance with the human proxy agent in all five evaluation layouts by setting the return as the utility function instead of regret.
In this experiment, return is a more appropriate utility function than regret in Human-AI coordination settings. The regret utility is not suitable for curating challenging environments that do not coordinate well with the environment buffer. Regret quantifies the opportunity cost of not making the optimal decision without considering coordination. Therefore, it is difficult to evaluate the effectiveness of coordination performance using regret. However, return is appropriate for evaluating the effectiveness of coordination performance because it is derived from the common rewards that agents receive through their collaborative efforts to achieve a common goal.
Learning Curve. Figure 4(a) shows the learning curves of our method and MAESTRO during training. By comparing these learning curves, we evaluate which utility function is more effective for training the ego-agent in Human-AI coordination. MAESTRO shows a high episode reward in the early training stage due to overfitting easy difficulty environments but high fluctuation when it encounters relatively less seen environments. On the other hand, our method steadily increases episode rewards by sampling more difficult environments according to the agent’s coordination ability. Appendix D demonstrates how the buffer composition of each method player changes as training progresses. Figure 4(b) illustrates the coordination performance of the ego-agent with the human proxy agent on randomly selected layouts from the training set at each training step. Meanwhile, Figure 4(c) shows the evaluation graph of the ego-agent coordination performance with the human proxy agent in the evaluation layouts. The use of return as the utility function leads to an increase in episode reward, whereas regret does not. This suggests that ego-agents trained with a return-based approach achieve more effective coordination with the human proxy agent.
@@ 5.3 Human-AI coordination
Human Study Result. Additionally, we conducted a human study to evaluate coordination with real humans. We recruited 20 participants from the local student community and evaluated the proposed method and the baseline models with our participants. Figure 5 shows the average coordination performance of five evaluation layouts in the human study. Our method outperformed the baseline models on all layouts ($test_{0}$, $test_{1}$, $test_{2}$, $test_{3}$) except for the hardest layout($test_{4}$). These results suggest that our method was robust to the unseen environment and diverse human partners.
Human Survey Result. For each session of the experiment, the the participants ranked the models in terms of collaborativeness (the most collaborative partner) and human preference (their top preference). The scores of the human evaluations in Figure 5 (b) were converted to a scale of 1 to 4 based on each participant’s subjective ranking of the models. A higher score indicates a better evaluation. The plot shows that our method has higher collaborativeness and human preference compared to the baseline model. Participants rated the agent trained with our method as more cooperative and preferred them for coordination over baseline models. The collaborativeness and human preference scores for each layout are provided in Appendix E.
@@ 6 Conclusion
Summary. In this paper, we propose an Automatic Curriculum Design for zero-shot Human-AI coordination. Our method extends the multi-agent Unsupervised Environment Design (UED) approach to zero-shot Human-AI coordination, which trains the ego-agent to coordinate with human partners in unseen environments. We use return as a utility function, unlike previous multi-agent UED methods that use regret. Our method outperformed against other baselines when evaluated with a human proxy in an Overcooked AI. The ablation studies demonstrate the effectiveness of the proposed components. Finally, we had a real human experiment to evaluate the human-AI coordination performance of our method. The result of human-AI coordination outperforms our method compared to other baselines.
Limitation and Future Work. When using the return as a measure of the learning potential of the environment/co-player pair, the evaluation performance improved compared to the previous method regret. However, prioritizing environments with low returns leads to the issue of sampling environment/co-player pair that are too challenging for the agent to learn effectively. In the future, we plan to introduce an alternative metric to promote the sampling of learnable environment-co-player pairs for Human-AI coordination Rutherford et al. (2025). We will also use mutation to adjust the environment that is sampled repeatedly to prevent overfitting the specific environments Parker-Holder et al. (2022); Wang et al. (2020).
In the current Multi-agent UED, the co-player pool is constructed from the past frozen weights of trained ego agents. This approach mitigates the non-stationarity problem compared to the self-play approach, which the co-player policy is changed during the training. However, since it essentially plays with the past of itself, it fails to train diverse strategies. We plan to study about how to increase the diversity of the co-player pool in UED.
@@
Our implementation is based on PLR Jiang et al. [2021b] and the Overcooked environment Carroll et al. [2019]. In all experiments, we train the ego agent using the PPO method Schulman et al. [2017]. The details of the network architecture and hyperparameter choices are provided in Table 1 and Table 2, respectively. Additionally, we discuss our hyperparameter selections in Section 4. Table 3 presents the final hyperparameter choices for all methods.
@@
Conv2D + LeakyReLU 25 5$\times$5
Conv2D + LeakyReLU 25 3$\times$3
Conv2D + LeakyReLU 25 3$\times$3
@@
Episodes $N$ 375
Overcooked. To evaluate the improvement of zero-shot human-AI coordination performance in unseen environments, we experiment in the Overcooked environment, where players can cook and serve the food. The environment has interactive blocks such as onions, plates, pots, and outlets, wall blocks that agents cannot move through, and space blocks that agents can move through. To get a reward from the environment, each player picks up 3 onions (3 reward) and puts them in the pot (3 reward). After 20 seconds, the onion soup is ready, take a plate to serve the onion soup (5 reward). They then submit the dish to the outlets. Finally, all players get a common reward of 20. This reward shaping is based on a previous Overcooked benchmark paper Carroll et al. [2019].
@@
1:  while Size of layout buffer $<$ $M$ do
@@
3:     Randomly place $N$ blocks
4:     if Duplicate(Layout, $L$then
@@
9:     if Sovability(Layout) then
@@
Algorithm 2 shows the overall layout generation process. A layout generator is a heuristic-driven tool whose input consists of three parameters: The maximum size of the layout buffer $M$, the number of interactive blocks $N$, and the number of empty spaces $E$. The generator ensures the presence of a minimum of $E$ spaces and $N$ interactive blocks within each map. The location of the block in the layout is selected randomly. In this experiment, we set the parameters as $M$=6,000, $N$= $6\sim 9$, and $E$=14. During the layout generation process, the generator checks the Hamming distance between embedded arrays to eliminate any redundancy. In addition, the solvability of each layout is checked using the A* path-finding algorithm Hart et al. [1968], which checks whether or not the agent can reach each interactive block.
@@
Figure 6 represents the detailed results of experiments with AI-AI cross-play on the five evaluation layouts. This is a zero-shot coordination experiment in which different models are coordinated without direct encounters with others during the training phase. These scores are normalized relative to the highest reward obtained from that layout, with the maximum value set at 1 and the minimum value set to 0. our method performs better than the baselines in the other five evaluation layouts. The final mean graph shows the average value of the normalized scores for the five layouts. Our method also outperforms the baseline in self-play that each method plays itself and in cross-play that each method plays against different methods.
PLR and MAESTRO use a regret-based utility function, which is not suitable for cooperative environments, resulting in a worse performance than the Random method. In contrast, our method improves the lower bound of the performance by using a return-based utility function. This approach curates challenging environment/co-player pairs to the ego-agent. Our trained ego-agent generally performs well in easy-to-hard environments and is robust to unseen environments and co-players.
Appendix D Co-player’s buffer analysis.
(a) Difficulty categorized by reward per layouts
(b) Layout difficulty composition in the co-player’s buffer
(c) Changing the layout difficulty composition of buffers
Figure 7: (a) Distribution of rewards obtained by the trained model on the training layout, playing with a human proxy. Difficulty is categorized based on these reward distributions. (b) Difficulty distribution of layouts sampled during the training process for the Our/MAESTRO method. (c) Difficulty configuration of the buffer changes as training progresses.
@@
Table 4: Difficulty classification based on reward conditions
Define training layouts difficulty. To categorize the difficulty of the layouts, we obtained the average reward (50 runs) evaluated from the trained agent and the human proxy agent. Figure 7(a) shows the distribution of average rewards obtained from the 6,000 layouts used for training. We apply the criteria outlined in Table 4, categorizing the difficulty levels into five groups, from ’Very Hard’ to ’Very Easy’, based on the mean and standard deviation. According to this classification scheme, the layouts classified as ‘Medium’ are the most numerous, while those classified as ‘Very Hard’ or ‘Very Easy’ are comparatively rare.
Co-player’s buffer Composition. Figure 7(b) shows the cumulative composition of layout difficulties in the environment buffer of the co-player, which is first added to the co-player population during the training process. Both our method and MAESTRO prioritize layout(environment) sampling and store layouts in the co-player buffer of size 1,000 according to their own criteria. In our method, the percentage of ’Very Hard’ layouts in the co-player buffer increases as training progresses from start to finish. This indicates that our method samples the more difficult layouts as the agent’s performance improves, similar to curriculum learning. However, MAESTRO was likely to sample easier environments as training progress, so the percentage of ’easy’ layouts in the co-player buffer increased. Figure7(c) represents details of the changing layouts from the co-player’s buffer. These results support the idea that the return-based utility function is more appropriate than the regret-based utility function to assess learning potential in the coordination setting.
Appendix E Additional human study analysis.
Figure 8: Human survey results by test layouts Results from a survey of 20 human players on five evaluation layouts. In the layouts of $test_{0},test_{1},test_{2}$, and $test_{3}$, human players tended to prefer our model, but in the hardest layout of $test_{4}$, the human player had different preferences, and each model scored similarly.

### v2

@@ 3.3 Learning Potential Score Based on Return
Our method follows the replay distribution (Equation 1) when selecting an environment for the ego-agent from an individual co-player buffer. In this replay distribution, $P_{S}$ is the first priority component that assigns higher priority to environments with learning potential for training the ego-agent.
As discussed in Section 3.2, return is the appropriate scoring metric to build a prioritization distribution in human-AI coordination. It represents the coordination performance during the episode. Previous works Zhao et al. (2023); Lou et al. (2023) have demonstrated that return captures coordination performance and also indicates which environments require more training. If the replayed environment has a low return, it indicates that the ego-agent is hard to coordinate in that environment. The details of $S$ are as follows:
@@
Equation (3) is the expected return of the replayed environment with the sampled co-player during the episode. Here, $T$ represents the total number of timesteps in the episode, $t\in\{0,\dots,T\}$ denotes each timestep within the episode, and $i$ denotes the index of the replayed environment from the co-player’s buffer. This environment score is normalized using Equation (4). This prioritization approach adapts the temperature-controlled ranking of previous work Jiang et al. (2021b) to our coordination setting. We normalize the scores using a rank-based prioritization, which sorts the scores in the co-player buffer in ascending order, prioritizing low-return environments. $j$ denotes the index of environments in the co-player’s buffer. $\beta\in[0,1]$ is the temperature parameter that adjusts the influence of the distribution.
@@ 4 Experiments Setup
Overcooked AI Carroll et al. (2019) is a cooperative game played in a grid-world environment inspired by the kitchen. The goal is to cook as much food as possible in a limited time. Agents can get rewards by completing tasks sequentially, such as preparing ingredients, cooking, and serving dishes. Cooperation is essential in this environment, where success depends on effective coordination between players and their partners.
Baseline Model. We compare our method with three key baseline models related to UED: MAESTRO Samvelyan et al. (2023), Robust PLR Jiang et al. (2021b), and Domain Randomization Jakobi (1997). Robust PLR and domain randomization are methods originally used in a single-agent setting, so we modified them for multi-agent settings using the self-play method. We used the PPO method Schulman et al. (2017) to train the ego-agent. PPO is the standard algorithm for training ego-agents through population-based learning with self-play in zero-shot human-AI coordination research. Previous studies Strouse et al. (2021) have demonstrated that this approach enables agents to coordinate with humans without relying on human data. The model architecture and hyperparameter choices are provided in Appendix A.
Layout Generation. In our experiments, we automatically generated a total of 6,000 different layouts for Overcooked-AI, each layout was 7 × 5. These layouts were generated using a heuristic-driven layout generator capable of adjusting three layout parameters: the number of interactive blocks (such as pots, plates, outlets, and onions), the number of spaces, and the size of the layouts, which is fixed in this study. The generator was designed to ensure that each layout included at least 12 spaces and 5 interactive blocks. To avoid redundancy, the layout generator compared a layout array during generation. In addition, we check the solvability of the layout using the A* path-finding algorithm Hart et al. (1968), which can find a path for the agent to access any interactive block. Further details on the algorithm are provided in the Appendix B.
Human Proxy. To evaluate the coordination performance of the trained agent in our experiments, we used a human proxy agent from a previous study Yang et al. (2022) as a co-player agent. This human proxy agent performs as human-like as possible by adjusting its parameters based on human data using cross-entropy methods.
Evaluation. We evaluated the human-AI zero-shot coordination performance on five unseen layouts (Figure 2), selected based on coordination difficulty. Specifically, the five layouts correspond to the 100th, 75th, 50th, 25th, and 0th percentiles of the maximum achievable self-play rewards, as in a previous study Yang et al. (2022). This helps to measure the adaptability of the ego-agent based on the difficulty of the layout. In addition, we conducted a utility function analysis.
@@
Figure 3: (a) Overall results with a human proxy model. Our method achieves higher coordination performance with the human proxy agent than other baselines on all five evaluation layouts (Mean and standard error were computed over 100 runs with three random seeds). (b) Utility function analysis (RegretRegret vs ReturnReturn). Using a return-based utility function has an advantage over a regret-based utility function in training an ego-agent for the regret-based utility function for multi-agent coordination settings (Mean and standard error were computed over 100 runs with three random seeds).
@@
Figure 4: Learning curves for the training, training-evaluation, and evaluation phases. The solid line shows the average episode reward across the three different random seeds, while the shaded areas show the maximum and minimum rewards. The results show that our model achieves higher performance on both the training and evaluation layouts.
Figure 5: Results of the prioritized co-player sampling Prioritized co-player sampling leads to higher coordination performance with the human proxy agent compared to random co-player sampling across all five evaluation layouts. (Mean and standard error were computed over 100 runs with three random seeds.) OursR represents our proposed method with random co-player sampling.
(a) Human Study Result.
(b) Human Survey Result.
Figure 6: This plot shows the results of an experiment with 20 real humans: (a) The average reward obtained on the five unseen evaluation layouts, (b) The average evaluation scores for the two items "Collaborativeness" and "Human Preference". Each subject was scored on a scale of 1 to 4 based on their voting ranking. A higher score indicates a more positive evaluation.
@@ 5.1 Experiments With Human Proxy
Overall Result. Figure 3(a) shows the comparison result of our method and the baseline models (MAESTRO, Robust PLR, and Domain Randomization) in terms of coordination performance with a human proxy agent on unseen layouts. The bar graph shows the evaluation results, which are the average episode rewards (three random seeds) with the human proxy agent on the evaluation layouts. As a result, our method showed better performance in all five evaluation layouts compared to other baselines. This performance gap is statistically significant across all layouts except $test_{4}$, as shown by the results of the paired t-test in Appendix E (see Table 5). We observe that our method achieves a higher generalization coordination performance with unseen partners on unseen evaluation layouts. We also performed cross-play experiments, where our method outperformed other baselines in unseen evaluation layouts. The results of these experiments are provided in the Appendix C.
@@ 5.2 Study of the Utility Function
Evaluation Result. We conducted an analysis to try to figure out which utility functions are appropriate for a human-AI coordination setting. Figure 3(b) shows how the utility function affects the training of the ego-agent. Our method and Robust PLR (return) outperformed MAESTRO (regret-based baseline) and Robust PLR in all five evaluation layouts with the human proxy agent, demonstrating the effectiveness of return over regret as a utility function.
In this experiment, return is a more appropriate utility function than regret in human-AI coordination settings. The regret utility function is not suitable for curating challenging environments for coordination tasks. Regret quantifies the opportunity cost of not making the optimal decision without considering coordination. Therefore, it is difficult to assess the effectiveness of coordination performance using regret. However, return is appropriate for evaluating coordination performance because it is derived from the common rewards that agents receive through their collaborative efforts to achieve a common goal.
Learning Curve. Figure 4(a) shows the learning curves of our method and MAESTRO during training. By comparing these learning curves, we evaluate which utility function is more effective for training the ego-agent in human-AI coordination. MAESTRO shows high episode reward in the early training stage due to overfitting easy difficulty environments, but shows high fluctuation when it encounters relatively less seen environments. On the other hand, our method steadily increases episode rewards by sampling more difficult environments according to the agent’s coordination ability. Appendix D shows how the buffer composition of each method changes as the training progresses. Figure 4(b) illustrates the coordination performance of the ego-agent with the human proxy agent on randomly selected layouts from the training set at each training step. Meanwhile, Figure 4(c) shows the evaluation graph of the coordination performance of the ego agent with the human proxy agent in the evaluation layouts. The use of return as a utility function leads to an increase in episode rewards, whereas regret does not. This suggests that the ego-agent trained with a return-based approach achieves more effective coordination with the human proxy agent.
5.3 Study of the Co-player Sampling
Evaluation Result. To evaluate the effect of prioritized co-player sampling, we compared our proposed method with random co-player sampling. Figure 5 shows that prioritized co-player sampling improves the training performance of the ego-agent. Our sampling strategy prioritizes co-players with low returns and jointly considers the challenging environments they have encountered. As a result, the ego-agent optimizes the lower bound of coordination performance. In contrast, random co-player sampling does not account for the varying difficulty levels of co-players or environments. This comparison highlights the importance of curating challenging coordination scenarios for the ego-agent.
@@ 5.4 Human-AI coordination
Human Study Result. Additionally, we conducted a human study to evaluate coordination with real humans. We recruited 20 participants from the local student community and evaluated the proposed method and the baseline models with our participants. All participants provided their informed consent before the experiment, which was conducted following ethical guidelines approved by the Institutional Review Board (IRB).
Figure 6 shows the average coordination performance of five evaluation layouts in the human study. Our method outperformed the baseline models on all layouts ($test_{0}$, $test_{1}$, $test_{2}$, $test_{3}$), but showed a similar performance to other baselines in $test_{4}$, with no statistically significant differences ($p>0.05$) according to a t-test. This performance is a limitation of our approach. Our method is designed to improve the lower bound of coordination performance by prioritizing training with difficult environment/co-player pairs, which enhances robustness in most scenarios. However, this strategy may bias the model towards specific coordination patterns, potentially limiting generalization in extremely complex environments like $test_{4}$. The sophisticated coordination requirements of $test_{4}$ may fall outside the range of patterns for which our method has been optimized. However, consistent performance improvements in $test_{0}$ through $test_{3}$ support the effectiveness of our method. The detailed results of the paired t-tests for all layouts are provided in Appendix E (see Table 6).
Human Survey Result. For each session of the experiment, the participants ranked the models in terms of collaborativeness (the most collaborative partner) and human preference (their top preference). The human evaluation scores in Figure 6 (b) were converted to a scale of 1 to 4 based on each participant’s subjective ranking of the models. A higher score indicates a better evaluation. The plot shows that our method has higher collaborativeness and human preference compared to the baseline models. Participants rated the agent trained with our method as more cooperative and preferred it for coordination over baseline models. Collaborativeness and human preference scores for each layout are provided in the appendix D.
@@ 6 Conclusion
Summary. In this paper, we propose Automatic Curriculum Design for Zero-Shot Human-AI coordination. Our method extends the multi-agent Unsupervised Environment Design (UED) approach to zero-shot human-AI coordination, which trains the ego-agent to coordinate with human partners in unseen environments. We use return as a utility function, unlike previous multi-agent UED methods that use regret. Our method outperformed other baselines when evaluated with a human proxy in Overcooked-AI. Ablation studies demonstrate the effectiveness of the proposed components. Finally, we had a real human experiment to evaluate the human-AI coordination performance of our method. The result of the human-AI coordination experiment shows that our method outperforms other baselines.
Limitation and Future Work. When using the return as a measure of the learning potential of the environment/co-player pair, the evaluation performance improved compared to a previous method. However, prioritizing environments with low returns leads to the issue of sampling environment/co-player pairs that may be too challenging for the agent to learn effectively. In future work, we plan to adopt an alternative metric based on success probability Rutherford et al. (2025) to promote the sampling of pairs of learnable environment-co-players for human-AI coordination. The success of coordination can be evaluated through indicators such as collision-free movements and cooperative completion of cooking tasks. We will also use mutation to adjust environments that are sampled repeatedly to prevent overfitting to specific environments Parker-Holder et al. (2022); Wang et al. (2020).
In the current multi-agent UED, the co-player pool is constructed from the past frozen weights of trained ego-agents. This approach mitigates the non-stationarity problem compared to self-play approaches, in which the co-player policy is changed during training. However, since it essentially plays with past versions of itself, it struggles to train diverse strategies. We plan to investigate methods for increasing the diversity of the co-player pool in UED.
@@
Our implementation is based on PLR Jiang et al. [2021b] and Overcooked-AI Carroll et al. [2019]. In all experiments, we train the ego-agent using the PPO method Schulman et al. [2017]. The details of the network architecture and hyperparameter choices are provided in Tables 1 and Table 2, respectively. Additionally, Table 3 presents the final hyperparameter choices for all methods.
@@
Conv2D + LeakyReLU 25 $\times5$
Conv2D + LeakyReLU 25 $\times3$
Conv2D + LeakyReLU 25 $\times3$
@@
Overcooked-AI. To evaluate the improvement in zero-shot human AI coordination performance in unseen environments, we experiment in the Overcooked AI environment, where players can cook and serve food. The environment has interactive blocks such as onions, plates, pots, and outlets, wall blocks that agents cannot move through, and space blocks that agents can move through. To get a reward from the environment, each player picks up 3 onions (3 reward) and puts them in the pot (3 reward). After 20 seconds, the onion soup is ready. Take a plate to serve the onion soup (5 reward). They then submit the dish to the outlets. Finally, all players get a common reward of 20. This reward shaping is based on a previous Overcooked-AI benchmark paper Carroll et al. [2019].
@@
Episodes NN 375
@@
1:while Size of layout buffer << $M$ do
@@
3:  Randomly place NN blocks
4:  if Duplicate(Layout, LLthen
@@
9:  if Solvability(Layout) then
@@
The algorithm 2 shows the overall layout generation process. A layout generator is a heuristic-driven tool whose input consists of three parameters: The maximum size of the layout buffer $M$, the number of interactive blocks $N$, and the number of empty spaces $E$. The generator ensures the presence of a minimum of $E$ spaces and $N$ interactive blocks within each map. The locations of the blocks in the layout are randomly selected. In this experiment, we set the parameters as $M=6,000$, $N=6\sim 9$, and $E=14$. During the layout generation process, the generator checks the Hamming distance between the embedded arrays to eliminate any redundancy. Additionally, we verify the solvability of each layout using the A* path-finding algorithm Hart et al. [1968], which determines whether the agent can reach each interactive block.
@@
Figure 7 represents the detailed results of experiments with AI-AI cross-play on the five evaluation layouts. This is a zero-shot coordination experiment in which different models are coordinated without direct encounters with others during the training phase. These scores are normalized relative to the highest reward obtained from that layout, with the maximum value set at 1 and the minimum value set to 0. Our method performs better than the baselines in all five evaluation layouts. The final mean graph shows the average value of the normalized scores for the five layouts. Our method also outperforms baselines in self-play where each method plays itself, and in cross-play where each method plays against different methods.
Robust PLR and MAESTRO use a regret-based utility function, which is not suitable for cooperative environments, resulting in worse performance than the random method. In contrast, our method improves the lower bound of the performance by using a return-based utility function. This approach curates challenging environment/co-player pairs for the ego-agent. Our trained ego-agent generally performs well in easy-to-hard environments and is robust to unseen environments and co-players.
Appendix D ADDITIONAL RESULTS.
Table 4: Difficulty classification based on reward conditions
@@
Define training layouts difficulty. To categorize the difficulty of the layouts, we obtained the average reward (50 runs) evaluated from the trained agent and the human proxy agent. Figure 9(a) shows the distribution of average rewards obtained from the 6,000 layouts used for training. We applied the criteria outlined in Table 4, categorizing the difficulty levels into five groups, from ’Very Easy’ to ’Very Hard’, based on the mean and standard deviation of the reward distribution. According to this classification scheme, the layouts classified as ’Medium’ are the most numerous, while those classified as ’Very Hard’ or ’Very Easy’ are comparatively rare.
Co-player’s buffer Composition. Figure 9(b) shows the cumulative composition of layout difficulties in the environment buffer of the co-player, which is first added to the co-player population during the training process. Both our method and MAESTRO prioritize layout(environment) sampling and store layouts in the co-player’s buffer of size 1,000 according to their own criteria. In our method, the percentage of ’Very Hard’ layouts in the co-player’s buffer increases as the training progresses from start to finish. This indicates that our method samples the more difficult layouts as the agent’s performance improves, similar to curriculum learning. However, MAESTRO was likely to sample easier environments as training progressed, so the percentage of ’Easy’ layouts in the co-player buffer increased. Figure 9(c) represents details of the changing layout composition in the co-player’s buffer. These results support the idea that the return-based utility function is more appropriate than the regret-based utility function to assess learning potential in a coordination setting.
(a) Difficulty categorized by reward per layouts
(b) Layout difficulty composition in the co-player’s buffer
(c) Changing the layout difficulty composition of buffers
Figure 8: (a) Distribution of rewards obtained by the trained model on training layouts, playing with a human proxy. Difficulty is categorized based on these reward distributions. (b) Difficulty distribution of layouts sampled during the training process for our method and MAESTRO. (c) The difficulty configuration of the buffer changes as training progresses.
Figure 9: Human survey results by test layouts Results from a survey of 20 human players on five evaluation layouts. In the layouts of $test_{0},test_{1},test_{2}$ and $test_{3}$, human players tended to prefer our model, but in the hardest layout of $test_{4}$, the human player had different preferences, and each model scored similarly.
Appendix E ADDITIONAL Paired t-test.
Table 5: Paired t-test results between Ours and baseline methods on each evaluation layout with a human proxy.
Layout Comparison t-statistic p-value
$test_{0}$ MAESTRO 4.8088 0.0406
Robust PLR 4.4569 0.0468
Random 4.9371 0.0387
$test_{1}$ MAESTRO 29.0707 0.0012∗∗
Robust PLR 9.4667 0.0110
Random 10.1263 0.0096∗∗
$test_{2}$ MAESTRO 7.6474 0.0167
Robust PLR 207.0434 <<0.0001∗∗∗
Random 25.8937 0.0015∗∗
$test_{3}$ MAESTRO 17.2566 0.0033∗∗
Robust PLR 10.8919 0.0083∗∗
Random 4.1389 0.0537
$test_{4}$ MAESTRO 3.5168 0.0722
Robust PLR 4.0035 0.0571
Random 1.7264 0.2264
Table 6: Paired t-test results between Ours and baseline methods on each layout evaluated with a human partner.
Layout Method t-statistic p-value
$test_{0}$ MAESTRO 3.2151 0.0046∗∗
Robust_PLR 3.6306 0.0018∗∗
Random 2.9939 0.0075∗∗
$test_{1}$ MAESTRO 6.1094 <0.0001∗∗∗
Robust_PLR 3.3245 0.0036∗∗
Random 0.9126 0.3729
$test_{2}$ MAESTRO 4.8766 0.0001∗∗∗
Robust_PLR 5.6865 <0.0001∗∗∗
Random 3.9149 0.0009∗∗∗
$test_{3}$ MAESTRO 4.3184 0.0004∗∗∗
Robust_PLR 3.3836 0.0031∗∗
Random 4.2080 0.0005∗∗∗
$test_{4}$ MAESTRO -0.4229 0.6771
Robust_PLR 0.5977 0.5571
Random -1.6296 0.1197